OpenStudy (anonymous):

fan and medal awarded m(m - 3) = 54

OpenStudy (cggurumanjunath):

open the bracket and multiply tell me what u get ?

OpenStudy (anonymous):

open the bracket ? what ???

OpenStudy (anonymous):

distribute m

OpenStudy (anonymous):

Ok I need the steps to be explained can anyone do that in plain English ??

OpenStudy (cggurumanjunath):

m(m-3)=54 m^2-3m=54 now solve for m

OpenStudy (anonymous):

still confused.

OpenStudy (cggurumanjunath):

\[m(m-3)=54\]

OpenStudy (amtran_bus):

m^2 -3m=54

OpenStudy (cggurumanjunath):

\[m^{2}-3m=54\]

OpenStudy (amtran_bus):

factor

OpenStudy (cggurumanjunath):

now factorise to find the value of m .

OpenStudy (cggurumanjunath):

do u know how to factorise ?

OpenStudy (cggurumanjunath):

@mikaa_toxica13

OpenStudy (cggurumanjunath):

m2−3m-54=0

OpenStudy (anonymous):

ok so m^2 would turn into 2m ?

OpenStudy (cggurumanjunath):

\[m ^{2}−3m-54=0\]

OpenStudy (anonymous):

yeaa I saw that. so it would turn into 2m ?

OpenStudy (cggurumanjunath):

no the previous was a mistake it is not m2−3m-54=0 it is m^2−3m−54=0

OpenStudy (anonymous):

yes so after I multiply the m^2 it would be 2m ?????????????????????

OpenStudy (cggurumanjunath):